(Clear, Alaska) – Special Notes: Today will be the last day Northwest Team 10 has command of the Clear Fire. At 6:00 p.m. command will transition to Northern Rockies Team 1. Members of both teams worked together yesterday to ensure a smooth transition and will do so again through 6:00 p.m. today. Northwest Team 10 Incident Commander Al Lawson wishes to thank our cooperators, stakeholders and local communities, collectively “Team Clear”, for their cooperation and for the confidence placed in us during our tenure at the Clear Fire. It was an honor to serve this community.
The Teklanika and Nenana Rivers have crested and are not expected to rise further, therefore, the recent flood watch has been canceled.
Weather: A cold front moves in today from the southwest with temperatures of 64F to 69F and humidity of 55-60%. Showers will move into the area around 3:00 p.m. and continue throughout the entire day. Winds should be from the southwest at 6-10 mph with gusts to 18 mph. Winds near the Parks Highway gap should be south to southwest at the same speeds. Additional moisture is forecast for Saturday. Combined with today’s precipitation, this should result in rainfall amounts up to 0.50 inches. Sunday and Monday bring more showers to the area, adding an additional 0.25 to 0.50 inches.
Fire Activity: Containment of the Clear Fire has reached 56%. There was drying and solar heating yesterday, and this did not prompt any significant increases in fire activity. Fire behavior continues to be mostly smoldering and creeping with the occasional isolated single tree torch. Aircraft identified some smoking areas in Div O, but these are well within the interior of the burned area. Fire behavior analysts anticipate that fire activity will remain mild in the coming days as more moisture moves into the region.
In Div K crews are making progress on fire line along the south-east most perimeter, south of the Kobe Rd. They are working to tie this line into a body of water and were near this goal by the end of yesterday. Crews will keep working this line today and will continue adding width to existing fire lines. Firefighters in Kobe Ag are conducting road repair with graders and have made significant progress on several roads. Masticator operators are improving conditions by removing and chipping fuels. The Structure Group remains in Div K locating areas of heat near structures and mopping them up when found. They are also ensuring structure protection equipment is maintained by testing pumps and sprinkler systems.
Yesterday, firefighters in Div G on the north side of Kobe Ag helped transport formerly spiked out crews back to the Incident Command Post. Crews are continuing backhaul and are doing suppression repair work in Div G.
Crews from Div B on the northeast assisted with helicopter shuttles yesterday to move crews from the fire area. Div B remains in patrol status.
Div A/O on the north and northwest is in patrol status, and crews remain in the area to act on any needs.
Sheltering or evacuation-related information: Contact the Denali Borough at (907) 683-1330, see the Borough’s Facebook page or visit denaliborough.org Fire loss or suppression repair information: Contact the Denali Borough at (907) 683-1330.
All other fire information: Clear Fire Information, (907) 290-2921; Alaska Fire Information, (907) 356-5511; akfireinfo.com, inciweb.nwcg.gov/incident/8178, and on the Alaska Division of Forestry and BLM Alaska Fire Service Facebook pages.